Use the table below to answer the questions that follow
X| 0 | 1| 2 | 3 | 4| 5|
Freq| 15| 24| 8| 6| 10| 4|

i) Use the table above to construct cumulative frequency and relative cumulative frequencies for the distribution
ii) Estimate the mean deviation
1
Expert's answer
2021-07-19T05:45:42-0400

Solution:

"Mean=\mu=\dfrac{\Sigma f_ix_i}{\Sigma f_i}
\\=\dfrac{15\times 0+24\times 1+8\times 2+6\times 3+10\times 4+4\times 5}{67}"

"=\dfrac{0+24+16+18+40+20}{67}
\\=\dfrac{118}{67}=1.76"



Now, mean deviation

"=\dfrac{\Sigma|x_i -\mu|}{\Sigma f_i}
\\=\dfrac{1.76+0.76+0.24+1.24+2.24+3.24}{67}
\\=0.14149"